OpenStudy (anonymous):

find the value of x if a line having a slope of 4 and passing through point (3,8) intersects the x-axis.

OpenStudy (anonymous):

@Xeph

OpenStudy (anonymous):

what's the equation of the line?

OpenStudy (anonymous):

I'm going to bed now but I'll give you a few tips that should get you through this problem

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

First, find the equation of your line by using point-slope form (which you just started using :D)

OpenStudy (anonymous):

can i use this formula y-y1=m(x-x1)

OpenStudy (anonymous):

That is point-slope form of a line. you use that to find your line's equation in this type of problem

OpenStudy (anonymous):

since the question gives you a point, and it gives you a slope

OpenStudy (anonymous):

After you have your line's equation, consider when the line crosses the x-axis (you're looking for x at that point)

OpenStudy (anonymous):

can i use this formula y-y1=m(x-x1)

OpenStudy (anonymous):

yes or no

OpenStudy (anonymous):

here's a hint: think about what y is equal to, at the point where the line crosses the x axis. then solve for x

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

I already answered that question. That is the point-slope equation of a line. If you want the equation for YOUR line, then substitute into that formula to obtain it

OpenStudy (anonymous):

when you have your line's equation, you can solve for x or y at various points on your line

OpenStudy (anonymous):

ok thank you :)

OpenStudy (anonymous):

so i got thr answer x=1

OpenStudy (anonymous):

you're welcome! good night, good luck on the rest of your math :)

OpenStudy (anonymous):



OpenStudy (anonymous):

thats the answer?

OpenStudy (anonymous):

@Xeph